mathslover (mathslover):

Resolve into factors : 99899

OpenStudy (anonymous):

Can u plz elaborate

mathslover (mathslover):

only this is given to us

OpenStudy (anonymous):

Like: 10=2*5

mathslover (mathslover):

this is a very very TRICKY quest. yes like that

hartnn (hartnn):

283*353

hartnn (hartnn):

but no idea how.

mathslover (mathslover):

http://www.meritnation.com/discuss/question/1050952

mathslover (mathslover):

answer is correct ... but... soln is ... not given ... try it

hartnn (hartnn):

oh wait, i m getting something....

hartnn (hartnn):

100000-101

mathslover (mathslover):

no

OpenStudy (anonymous):

I am sure it has something to do with a^2-b^2

mathslover (mathslover):

try but no..

OpenStudy (experimentx):

try this (100a+10b+c)(100d+10e+f) = 99899

mathslover (mathslover):

should I give out the soln ? tht I tried (help of vishwesh)

OpenStudy (anonymous):

wait a minute

OpenStudy (anonymous):

FACTORISE: 9x^4 +9x^3+ 8x^2+9x+9

OpenStudy (anonymous):

x=10

OpenStudy (anonymous):

So, that was right

mathslover (mathslover):

makes sense? yes saurav

OpenStudy (anonymous):

Yes

OpenStudy (experimentx):

pretty elegant technique.

OpenStudy (anonymous):

*

mathslover (mathslover):

yep it was sample paper of JMO

OpenStudy (experimentx):

what's JMO?

mathslover (mathslover):

http://www.icbse.com/papers/kvs-jmo-question-paper

OpenStudy (anonymous):

I have heard this olympiad word many times in OS........ actually what is meant by this?

OpenStudy (anonymous):

@mathslover can u plz explain what is meant by it?

OpenStudy (anonymous):

Is it a competition ?

mathslover (mathslover):

exam after which the selected in this exam get training from teachers for INMO

mathslover (mathslover):

INDIAN NATIONAL MATHIS OLYMPIAD

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

@experimentX do we have that in our country?

OpenStudy (anonymous):

Never heard of such things

OpenStudy (experimentx):

nope ... though you can still participate at IMO

mathslover (mathslover):

The four digit number aabb is a square number. Find it. --> Eg.

mathslover (mathslover):

If the number = N N = 1000x + 100x + 10y + y = 11 (100x + y) N is a perfect square divisible by 11. So, 100x + y is divisible by 11.  100x + y = 9 x 11x + x + y so x + y is divisible by 11. x and y < 9 and x and y  0 so, 1 < x + y < 18 so x + y = 11  N = 11 (100x + y) = 11 {99x + (x+y)} = 11 (99x + 11) = (11)2 (9x + 1) so 9x + y must be a perfect square and 1 < x < 9 so 9x + y is perfect square when x = 7  N = (11)2 x 64 = 7744

mathslover (mathslover):

copied ...

OpenStudy (experimentx):

@sauravshakya try contacting your school ... you might qualify.

hartnn (hartnn):

i wish i have heard of it when i was of your age.....

OpenStudy (anonymous):

Thanx @experimentX ........ But I have no idea about it.

OpenStudy (experimentx):

try contacting your school ... Olympiad problem are pretty difficult. I can't solve most of them. mukushla also get's quite trouble with some of them. probably they set of some test to qualify for it. I bet no one from this has ever won.

OpenStudy (anonymous):

If u two cant then how can I.....

OpenStudy (experimentx):

you can't go once you are after out of school ... also there is physics olympiad. these guys who win math olympiad become field medalist.

OpenStudy (experimentx):

*in the future. keep trying ... i (might) have tried if had enough time ... I can't waste my time on these days. probably i might be working from next year.

OpenStudy (experimentx):

my golden days are gone :(

OpenStudy (anonymous):

I think your golden days are ahead of u

OpenStudy (anonymous):

I know u and mukushla has got very high skill

OpenStudy (experimentx):

not really ... i rely mostly on experience.

OpenStudy (anonymous):

i heard that @mathslover said he cant take part in olympiads...why ??

OpenStudy (anonymous):

u cant go to ur country national team??

OpenStudy (mayankdevnani):

@mathslover prime factors?????

OpenStudy (mayankdevnani):

The factors of 99899 are 1, 283, 353 and 99899.

OpenStudy (mayankdevnani):

@mathslover reply.....plz....

OpenStudy (mayankdevnani):

coz 99899 = 283 * 353

mathslover (mathslover):

@mukushla we can try to go next year not now :)